# Solar Neutrinos and the Strongest Oscillation Constraints on Scalar NSI

### v1

@@ 5 Solar data analysis in presence of non-diagonal scalar NSI parameters
Figure 8: 90% C.L. contours in the $|\eta_{\alpha\beta}^{\odot}|$ vs $\phi_{\alpha\beta}$ planes for the three non-diagonal scalar NSI parameters.
@@ 6 The effect of the priors on the solar parameters
Regarding JUNO, we expect that that the prior of the mass splitting will only play a minor role with respect to the KamLAND case. Indeed, given that solar data can only weakly constrain the mass splitting, both priors have the effect of fixing the mass splitting in our analyses. On the other hand, the strong bound on the mixing angle can have a strong impact on the allowed regions, reducing them in the case of $\eta_{\mu\mu}^{\odot}$, $\eta_{e\mu}^{\odot}$ and $\eta_{\mu\tau}^{\odot}$. In the case of $\eta_{ee}^{\odot}$, $\eta_{\tau\tau}^{\odot}$ and $\eta_{e\tau}^{\odot}$, on the other hand, we observe a more pronounced shape change in the allowed regions. This is made possible by the fact that the preferred value of $\theta_{12}$ using solar data is larger than the one used to add the JUNO prior. This resulted in a slightly increased minimum $\chi^{2}$ at the $\lesssim 2\sigma$ level.
@@
All of these results are summarized in table 3, where we show the 90% CL limits on the parameters (1 degree of freedom). With $\Re\eta_{\alpha\beta}^{\odot}$ ($\Im\eta_{\alpha\beta}^{\odot}$) we refer to the limits on off-diagonal scalar NSI parameters obtained for $\phi_{\alpha\beta}=0,\pi$ ($\phi_{\alpha\beta}=\pm\pi/2$). We note that for comparison the Borexino only results from Ge:2018uhz translated to our units at 90% are approximately $-38\leq\eta_{ee}^{\odot}\leq 13$ and $-18\leq\eta_{\mu\mu,\tau\tau}^{\odot}\leq 27$ with a low significance preferred value of $\eta_{ee}^{\odot}=-36$. Thus our results with Solar and KamLAND data represent an improvement over existing constraints by about 1 to 1.5 orders of magnitude with the possibility for additional improvement with JUNO’s reactor measurement of the solar parameters. In the units of Babu:2019iml , we find that the upper limit on $\Delta m_{\rm Sun}\lesssim 0.8$ meV depending on the neutrino flavor and matter fermion, also an order of magnitude improvement over existing constraints. We note that, as in the case of Babu:2019iml , the constraint on the couplings $y_{a}y_{\alpha\beta}$ will weaken for mediators lighter than $m_{\phi}\sim 10^{-14}$ eV.
@@ 7 Conclusions
Finally, we point out the crucial fact that sNSI is fundamentally intertwined with the absolute neutrino mass scale and a possible detection of sNSI would also provide new and about the about the mass scale that is very complementary to other searches from cosmology or end point searches.
@@ Acknowledgements.
PBD acknowledges support from the US Department of Energy under Grant Contract DE-SC0012704.

### v2

@@ 5 Solar data analysis in presence of non-diagonal scalar NSI parameters
Figure 8: 90% CL contours in the $|\eta_{\alpha\beta}^{\odot}|$ vs $\phi_{\alpha\beta}$ planes for the three non-diagonal scalar NSI parameters.
@@ 6 The effect of the priors on the solar parameters
Regarding JUNO, we expect that the prior of the mass splitting will only play a minor role with respect to the KamLAND case. Indeed, given that solar data can only weakly constrain the mass splitting, both priors have the effect of fixing the mass splitting in our analyses. On the other hand, the strong bound on the mixing angle can have a strong impact on the allowed regions, reducing them in the case of $\eta_{\mu\mu}^{\odot}$, $\eta_{e\mu}^{\odot}$ and $\eta_{\mu\tau}^{\odot}$. In the case of $\eta_{ee}^{\odot}$, $\eta_{\tau\tau}^{\odot}$ and $\eta_{e\tau}^{\odot}$, on the other hand, we observe a more pronounced shape change in the allowed regions. This is made possible by the fact that the preferred value of $\theta_{12}$ using solar data is larger than the one used to add the JUNO prior. This resulted in a slightly increased minimum $\chi^{2}$ at the $\lesssim 2\sigma$ level.
@@
All of these results are summarized in table 3, where we show the 90% CL limits on the parameters (1 degree of freedom) and fig. 13. With $\Re\eta_{\alpha\beta}^{\odot}$ ($\Im\eta_{\alpha\beta}^{\odot}$) we refer to the limits on off-diagonal scalar NSI parameters obtained for $\phi_{\alpha\beta}=0,\pi$ ($\phi_{\alpha\beta}=\pm\pi/2$). We note that for comparison the Borexino only results from Ge:2018uhz translated to our units at 90% are approximately $-38\leq\eta_{ee}^{\odot}\leq 13$ and $-18\leq\eta_{\mu\mu,\tau\tau}^{\odot}\leq 27$ with a low significance preferred value of $\eta_{ee}^{\odot}=-36$. Thus our results with Solar and KamLAND data represent an improvement over existing constraints by about 1 to 1.5 orders of magnitude with the possibility for additional improvement with JUNO’s reactor measurement of the solar parameters. In the units of Babu:2019iml , we find that the upper limit on $\Delta m_{\rm Sun}\lesssim 0.8$ meV depending on the neutrino flavor and matter fermion, also an order of magnitude improvement over existing constraints. We note that, as in the case of Babu:2019iml , the constraint on the couplings $y_{a}y_{\alpha\beta}$ will weaken for mediators lighter than $m_{\phi}\sim 10^{-14}$ eV.
@@
Figure 13: The 1D $\Delta\chi^{2}$’s for the three diagonal parameters (top row) and the absolute value of the three off-diagonal parameters (bottom row). The solid lines are solar data only and the dashed lines are solar data and KamLAND. The dotted lines are the expected sensitivity including solar data, KamLAND, and JUNO. A horizontal line is drawn at the 90% CL point.
@@ 7 Conclusions
Finally, we point out the crucial fact that sNSI is fundamentally intertwined with the absolute neutrino mass scale and a possible detection of sNSI would also provide new information about the mass scale that is very complementary to other searches from cosmology or end point searches.
@@ Acknowledgements.
We thank Evgeny Akhmedov for helpful comments. PBD acknowledges support from the US Department of Energy under Grant Contract DE-SC0012704.